OpenStudy (anonymous):

Cylinder A has a radius of 1m and a height of 3m Cylinder B has a radius of 3 m and a height of 3 m what is the ration of volume of cylinder A to the volume of cylinder B?

iYuko (iyuko):

\[v=\Pi~r ^{2}h\]

iYuko (iyuko):

the ll is pi

iYuko (iyuko):

(a) v=3.14*1*1*3 v=9.42

iYuko (iyuko):

(b) v=3.14*3*3*3 v=84.78

OpenStudy (anonymous):

a. 1:1 b. 1:3 c. 1:9 d. 3:1 are my answer options @iyuko